# Extraordinary manifestation of near electrostatic field caused by macroscopic quantum shell effects in submicron hemispherical clusters

### v1

@@ Abstract
As shown in Ref. Brack (1993), Approach 1 is extensively applied for spherically symmetric clusters, as soon as the problem is reduced to the one-dimensional case. This allows to evaluate some properties of large clusters up to tens of thousands atoms. Approach 2 in general is related to quantum chemistry, allowing three-dimensional modeling of systems containing up to thousands of particles by means of modern supercomputing. Approach 3 may provide some insight to the properties of the system with rather small numerical cost and even allows to obtain compact analytical expressions by utilizing symmetries of the cluster.
Recently, Approach 3 was applied to the analysis of closed nonperiodic orbits in a spherical potential for describing oscillating contributions to the spatial distribution of electronic density Roccia and Brack (2008); Roccia et al. (2010). It was noted that there are two types of density oscillations: regular ones characterized by a small spatial scale, and irregular ones with a noticeably larger scale, which can exist only in spatial dimensions greater than one. However, there was quite small attention devoted to the analysis of spatial properties these oscillations, while the mathematical aspect of the problem was discussed in details.
@@
The effect is related to the manifestation of shell effects and can be intuitively understood using the semiclassical Green’s function method, which involves integrating over closed classical trajectories Gutzwiller (2013). The large-scale behavior is determined by the closed periodic trajectories. The major oscillations observed in the spatial distribution of electron density and electrostatic potential are determined by the first classical periodic trajectories of an electron, namely the $(3,1)$ and $(4,1)$ orbits in a spherically symmetric potential well Kuratov et al. (2019, 2021) as shown in Fig. 1a. Different possible effects related to macroscopic shell structure are considered in Refs. Kuratov et al. (2023, 2024). It has been shown that it leads to a modification of the spectrum of electronic oscillations in metal clusters, and also reveals a new mechanism of hydrodynamic cumulation initiated by quantum shell effects.
@@
Initial calculations are performed for a hemispherical cluster composed of $288$ Li atoms placed in the center of a $7\times 7\times 7$ nm3 cubic box. The atoms were organized in a BCC structure with a lattice parameter of $3.49$corresponding to normal material density and the radius of the modeled hemisphere is $1.4$ nm. The specificity of our calculations is that quite dense grids in real space are required. Even the most RAM-rich computers used in this work are unable to carry out computations on fine grids denser than $900\times 900\times 900$. The results for a $1.4$ nm hemisphere are obtained using the grid of $700\times 700\times 700$ cells (which means the cell size is less than $0.1$).
Besides the above described calculation, spherical and hemispherical lithium clusters with radii of $2$ nm were considered, which contain a significantly larger numbers of atoms: $1600$ and $960$, respectively. To control the influence of boundary conditions we additionally set the edge length $L$ of the cubic domain in simulations for hemispherical clusters to $9$, $12$, and $16$ nm. It is worth noting that the choice of lithium is driven by the possibility of using single-electron PAW pseudopotential in GPAW, so that the number of electrons in calculations is equal to number of ions, which allows to evaluate large clusters at less computational cost in a relatively large simulation domain.
@@
Applying the approximate model (2)–(3) yields the following values for the surface density and electric field strength of a metallic cluster with radius $1$ nm and $n_{e}=1.8\cdot 10^{29}$ m-3 near a flat surface
@@
Thus, the calculated field on a flat surface is $E\sim 10^{7}\,$/cm, which is close in order of magnitude to the calculated value of $E\sim 2.0\cdot 10^{6}\,$/cm and decreases linearly for $r\leq R$. We suppose the difference between the calculated and theoretical values is explained by the fact that the theoretical model does not take into account the near-surface electron distribution spreading, which reduces the calculated electric field value. At distances greater than the cluster radius, the field decreases inversely proportional to the square of the distance, which also agrees with the results of theoretical analysis.
@@
One may notice, that the effect of boundary conditions is negligible for the cubic domain edges of $L=12\,$nm and larger, so that such clusters may be considered as isolated ones. The result (9), obtained from theoretical analysis, agrees with the numerical calculations presented in Fig. 3c: the field near the flat surface of the hemisphere exists on the scale of the radius and decreases in magnitude according to Eq. (4).
The performed theoretical and numerical analysis allows us to estimate the magnitude of the electrostatic field for large cluster sizes, for which simulations are not possible:
@@
The analyzed effect can find numerous and effective applications in various fields of science and technology.
@@
The results of our study allow for another approach for electrostatic tweezers construction. A system of hemispheres supporting permanent electric polarization can be used, for example, to create SERS substrates that exhibit pincer-like properties. Local charges on the surface of hemispheres serve as sources of an electric field with a large spatial gradient, which should lead to the attraction of polarized molecules into the region of most intense field. As a result, an island film of such hemispheres can facilitate the deposition of molecules from a solution.
@@
The strong enhancement of local electric fields around nanoparticles has significant implications for various optical processes, including Surface enhanced Raman spectroscopy (SERS) Sharma et al. (2012). SERS is a technique that exploits the enhanced local electric fields to amplify the Raman signal from molecules adsorbed on the nanoparticle surface. A system of hemispheres being a source of electrical polarization can be used to create SERS substrates that simultaneously exhibit pincers-like properties. Namely, local charges on the surface of the hemispheres create an electric field with a large spatial gradient, which should lead to the attraction of polarized molecules into the maximum of the field. As a result, an island film of such hemispheres can facilitate the deposition of molecules from solution.
@@
Yet another application may include coating a system of hemispheres with a semiconductor that exhibits electro-optical properties, such as indium tin oxide (ITO) Ma et al. (2015). Local areas around these hemispheres will demostrate different semiconductor properties as the effect of absorption in the optical range will occur. The hemispheres can be arranged in groups of sizes of an order corresponding to the visible radiation wavelength. For ITO material to obtain metallic properties, a voltage of about $3$ V is required Ma et al. (2015). Our DFT results show (Fig. 3c) that the voltage arising sphere, on the order of $0.1$ V and can change the optical properties of ITO.

### v2

@@ Abstract
As shown in Ref. Brack (1993), Approach 1 is extensively applied for spherically symmetric clusters, as soon as the problem is reduced to the one-dimensional case. This allows to evaluate some properties of large clusters up to tens of thousands of atoms. Approach 2 in general is related to quantum chemistry, allowing three-dimensional modeling of systems containing up to thousands of particles by means of modern supercomputing. Approach 3 may provide some insight to the properties of the system with rather small numerical cost and even allows to obtain compact analytical expressions by utilizing symmetries of the cluster.
Recently, Approach 3 was applied to the analysis of closed nonperiodic orbits in a spherically symmetric potential for describing oscillating contributions to the spatial distribution of electronic density Roccia and Brack (2008); Roccia et al. (2010). It was noted that there are two types of density oscillations: regular ones characterized by a small spatial scale, and irregular ones with a noticeably larger scale, which can arise in a problem with more than one spatial dimension. However, there was quite small attention paid to the analysis of spatial properties these oscillations, while the mathematical aspect of the problem was discussed in details.
@@
The phenomenon is related to the manifestation of shell effects and can be intuitively understood using the semiclassical Green’s function method, which involves integrating over closed classical trajectories Gutzwiller (2013). The large-scale behavior is determined by the closed periodic trajectories. The major oscillations observed in the spatial distribution of electron density and electrostatic potential are determined by the first classical periodic trajectories of an electron, namely the $(3,1)$ and $(4,1)$ orbits in a spherically symmetric potential well Kuratov et al. (2019, 2021) as shown in Fig. 1a. Different possible effects related to macroscopic shell structure are considered in Refs. Kuratov et al. (2023, 2024). It has been shown that it leads to a modification of the spectrum of electronic oscillations in metal clusters, and also reveals a new mechanism of hydrodynamic cumulation driven by quantum shell effects.
@@
Initial calculations are performed for a hemispherical cluster composed of $288$ Li atoms placed in the center of a $7\times 7\times 7$ nm3 cubic box. The atoms were organized in a body-centered cubic (bcc) structure with a lattice parameter of $3.49$corresponding to normal material density and the radius of the modeled hemisphere is $1.4$ nm. The specificity of our calculations is that quite dense grids in real space are required. Even the most RAM-rich computers used in this work are unable to carry out computations on fine grids denser than $900\times 900\times 900$. The results for a $1.4$ nm hemisphere are obtained using the grid of $700\times 700\times 700$ cells (which means the cell size is less than $0.1$).
Besides the above described calculation, spherical and hemispherical lithium clusters with radii of $2$ and $2.5$ nm were considered, which contain significantly larger numbers of atoms: $960$ and $1600$, respectively. To control the influence of boundary conditions we additionally set the edge length $L$ of the cubic domain in simulations for hemispherical clusters to $9$, $12$, and $16$ nm. It is worth noting that the choice of lithium is driven by the possibility of using single-electron PAW pseudopotential in GPAW, so that the number of electrons in calculations is equal to number of ions, which allows to evaluate large clusters at less computational cost in a relatively large simulation domain.
@@
Applying the approximate model Eqs. (2)–(3) yields the following values for the surface density and electric field strength of a metallic cluster with radius $1$ nm and $n_{e}=1.8\cdot 10^{29}$ m-3 near a flat surface
@@
Thus, the calculated field on a flat surface is $E\sim 10^{7}\,$/cm, which is close in order of magnitude to the calculated value of $E\sim 2.0\cdot 10^{6}\,$/cm and decreases linearly for $r\leq R$. We suppose that the difference between the calculated and theoretical values is explained by the fact that the theoretical model does not take into account the near-surface electron distribution spreading, which reduces the calculated electric field value. At distances greater than the cluster radius, the field decreases inversely proportional to the square of the distance, which also agrees with the results of theoretical analysis.
@@
One may notice, that the effect of boundary conditions is negligible for the cubic domain edges of $L=12\,$nm and larger, so that such clusters may be considered as isolated ones. The result (9), obtained from theoretical analysis, agrees with the numerical calculations presented in Fig. 3c: the field near the flat surface of the hemisphere acts on the scale of the radius and decreases in magnitude according to Eq. (4).
The performed theoretical and numerical analyses allow us to estimate the magnitude of the electrostatic field for large cluster sizes, for which simulations are not possible:
@@
The analyzed effect can find numerous and effective applications in various areas of science and technology.
@@
The results of our study allow for another approach for electrostatic tweezers construction, including Surface enhanced Raman spectroscopy (SERS) Sharma et al. (2012).
@@
The strong enhancement of local electric fields around nanoparticles has significant implications for various optical processes. SERS is a technique that exploits the enhanced local electric fields to amplify the Raman signal from molecules adsorbed on the nanoparticle surface. A system of hemispheres being a source of electrical polarization can be used to create SERS substrates that simultaneously exhibit pincers-like properties. Namely, local charges on the surface of the hemispheres create an electric field with a large spatial gradient, which should lead to the attraction of polarized molecules into the maximum of the field. As a result, an island film of such hemispheres can facilitate the deposition of molecules from solution.
@@
Yet another application may include coating a system of hemispheres with a semiconductor that exhibits electro-optical properties, such as indium tin oxide (ITO) Ma et al. (2015). Local areas around these hemispheres will demostrate different semiconductor properties as the effect of absorption in the optical range will occur. The hemispheres can be arranged in groups which sizes are of the order of the visible radiation wavelength. For ITO material to obtain metallic properties, a voltage of about $3$ V is required Ma et al. (2015). Our DFT results show (Fig. 3c) that the optical properties of ITO may be changed.